OpenStudy (anonymous):

The product of ten times Karen's age and thirty-two less than eight times Karen's age is zero. Find Karen's age.

OpenStudy (anonymous):

4 is the answer let karen's age be x ten times Karen’s age =10x thirty-two less than eight times Karen’s age = 8x-32 product = 0 (10x) (8x-32) = 0 10x = 0 / (8x-32) 10x = 0 x = 0 [not possible] [ or] product = 0 (10x) (8x-32) = 0 8x-32 = 0 / (10x) 8x-32 = 0 8x = 32 x = 4 the age of karen is 4 years Source(s): myself

OpenStudy (anonymous):

Thanks Gio that made it easier to understand

OpenStudy (jdoe0001):

source, yourself? you're 4? well, you're doing terrificly well for a 4 years old :P

OpenStudy (anonymous):

lol you welcome @Gunsy and yes i am doing well @jdoe0001 lmbo